# A Bernstein polynomial approach for the estimation of cumulative distribution functions in the presence of missing data

### v1

@@ A Taylor expansion of $1/\hat{p}(\boldsymbol{X}_{i})$ around $1/p(\boldsymbol{X}_{i})$ yields
$F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)=F_{Y_{i}\mid\boldsymbol{X}_{i}}(y)+f_{Y_{i}\mid\boldsymbol{X}_{i}}(y)(k/m-y)+\mathcal{O}(|k/m-y|^{2}),$ (7.20)
@@
$n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)b_{m,k}(y)\right)^{2}\right]$
$\quad=n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right]$
$\qquad-2n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\sum_{k=0}^{m}(k/m-y)b_{m,k}(y)\right]$
$\qquad+\mathcal{O}\left(\frac{n^{-1}}{\pi_{\min}}\left(\sum_{k=0}^{m}(k/m-y)b_{m,k}(y)\right)^{2}\right).$
The second term on the right-hand side is zero because $\sum_{k=0}^{m}(k-my)b_{m,k}(y)=0$. The error term is $\mathcal{O}(n^{-1}m^{-1})$ by Jensen’s inequality since
@@
Therefore,
$n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)b_{m,k}(y)\right)^{2}\right]$ $=n^{-1}\mathsf{E}\left[\frac{(1-\pi_{1}(\boldsymbol{X}_{1}))}{\pi_{1}(\boldsymbol{X}_{1})}\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right]$
$\quad+\mathcal{O}(n^{-1}m^{-1}).$
Putting back this expression in (7.19) yields the conclusion. ∎
@@
Let $y\in(0,1)$. We utilize the asymptotic representation (7.2) derived in the proof of Proposition 7:
@@
$B_{n,m}=n^{-1}\sum_{i=1}^{n}U_{i,m},\quad U_{i,m}=\frac{(\delta_{i}-\pi_{i}(\boldsymbol{X}_{i}))}{\pi_{i}(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)b_{m,k}(y).$ (7.21)
@@
The $W_{i,m}$’s are iid and centered, and $\mathsf{Var}(W_{1,m})\to\nu^{2}(y)>0$ as $n\to\infty$ by Proposition 7.
@@
$\lim_{n\to\infty}\frac{1}{\mathsf{Var}(W_{1,m})}\mathsf{E}\left[W_{1,m}^{2}\mathds{1}_{\{|W_{1,m}|^{2}>\varepsilon n\,\mathsf{Var}(W_{1,m})\}}\right]=0.$ (7.22)
By Assumption (A1), $\pi_{i}(\boldsymbol{X}_{i})\geq\pi_{\min}>0$. Using $\sum_{k=0}^{m}b_{m,k}(y)=1$ and $0\leq\mathcal{B}_{m}(F)(y)\leq 1$,
@@
so $|W_{1,m}|\leq 2\pi_{\min}^{-1}+1<\infty$. Since $\mathsf{Var}(W_{1,m})\to\nu^{2}(y)>0$, we have $\varepsilon n\,\mathsf{Var}(W_{1,m})\to\infty$, hence for all sufficiently large $n$, the indicator $\smash{\mathds{1}_{\{|W_{1,m}|^{2}>\varepsilon n\,\mathsf{Var}(W_{1,m})\}}}$ is equal to zero almost surely. Therefore the expectation in (7.22) is eventually 0, and the Lindeberg condition holds. ∎
@@ Appendix A Reproducibility
The R code that generated the figures, the simulation study results, and the real-data application is available online in the GitHub repository of Gharbi et al. , (2025).
@@ Appendix B List of acronyms
iid independent and identically distributed

### v2

@@ A Taylor expansion of $1/\hat{p}(\boldsymbol{X}_{i})$ around $1/p(\boldsymbol{X}_{i})$ yields
$F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)=F_{Y_{i}\mid\boldsymbol{X}_{i}}(y)+f_{Y_{i}\mid\boldsymbol{X}_{i}}(y)(k/m-y)+\mathcal{O}_{y}(|k/m-y|^{2}),$ (7.20)
@@
$n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)\,b_{m,k}(y)\right)^{2}\right]$ (7.21)
$\quad=n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right]$
$\qquad-2n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\sum_{k=0}^{m}(k/m-y)\,b_{m,k}(y)\right]$
$\qquad+\mathcal{O}_{y}\left(\frac{n^{-1}}{\pi_{\min}}\left(\sum_{k=0}^{m}(k/m-y)\,b_{m,k}(y)\right)^{2}\right).$
The second term on the right-hand side is zero because $\sum_{k=0}^{m}(k-my)\,b_{m,k}(y)=0$. The error term is $\mathcal{O}_{y}(n^{-1}m^{-1})$ by Jensen’s inequality since
@@
The error term depends on $y$ in (7.21) because $f_{Y_{1}\mid\boldsymbol{X}_{1}}^{\prime}$ may not be bounded on $(0,1)$. Therefore,
$n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\left(\sum_{k=0}^{m}F_{Y_{1}\mid\boldsymbol{X}_{1}}(k/m)\,b_{m,k}(y)\right)^{2}\right]$ $=n^{-1}\,\mathsf{E}\left[\frac{(1-\pi(\boldsymbol{X}_{1}))}{\pi(\boldsymbol{X}_{1})}\{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)\}^{2}\right]$
$\quad+\mathcal{O}_{y}(n^{-1}m^{-1}).$
Substituting this expression back into (7.19) yields the conclusion. ∎
@@
Let $y\in(0,1)$ be given. We utilize the asymptotic representation (7.2) derived in the proof of Proposition 7:
@@
$B_{n,m}=n^{-1}\sum_{i=1}^{n}U_{i,m},\qquad U_{i,m}=\frac{(\delta_{i}-\pi(\boldsymbol{X}_{i}))}{\pi(\boldsymbol{X}_{i})}\sum_{k=0}^{m}F_{Y_{i}\mid\boldsymbol{X}_{i}}(k/m)\,b_{m,k}(y).$ (7.22)
@@
The $W_{i,m}$’s are i.i.d. and centered, and $\mathsf{Var}(W_{1,m})\to\nu^{2}(y)>0$ as $n\to\infty$ by Proposition 7.
@@
$\lim_{n\to\infty}\frac{1}{\mathsf{Var}(W_{1,m})}\mathsf{E}\left[W_{1,m}^{2}\mathds{1}_{\{|W_{1,m}|^{2}\,>\,\varepsilon\hskip 0.56905ptn\,\mathsf{Var}(W_{1,m})\}}\right]=0.$ (7.23)
By Assumption (A1), $\pi(\boldsymbol{X}_{i})\geq\pi_{\min}>0$. Using $\sum_{k=0}^{m}b_{m,k}(y)=1$ and $0\leq\mathcal{B}_{m}(F)(y)\leq 1$,
@@
so $|W_{1,m}|\leq 2\pi_{\min}^{-1}+1<\infty$. Since $\mathsf{Var}(W_{1,m})\to\nu^{2}(y)>0$, we have $\varepsilon\hskip 0.56905ptn\,\mathsf{Var}(W_{1,m})\to\infty$, hence for all sufficiently large $n$, the indicator $\smash{\mathds{1}_{\{|W_{1,m}|^{2}\,>\,\varepsilon\hskip 0.56905ptn\,\mathsf{Var}(W_{1,m})\}}}$ is equal to zero almost surely. Therefore the expectation in (7.23) is eventually 0, and the Lindeberg condition holds. ∎
@@ Appendix A Reproducibility
The R code that generated the figures, the simulation study results, and the real-data application is available online in the GitHub repository of Gharbi et al. (2026).
@@ Appendix B List of acronyms
i.i.d. independent and identically distributed
I-IPW integrated inverse probability weighted
@@
Acknowledgments
The authors thank the three referees for their careful reading and constructive comments, which led to substantial improvements in the clarity and quality of the paper.
Funding
The work of Wissem Jedidi is supported by the Ongoing Research Funding program (ORF-2026-162) at King Saud University, Riyadh, Saudi Arabia. Frédéric Ouimet is supported by the start-up fund (1729971) from the Université du Québec à Trois-Rivières.